# Beyond Expected Goals: A Probabilistic Framework for Shot Occurrences in Soccer

### v1

@@ 1.1 Expected Goals
Expected Goals (xG) has become the most commonly used metric in modern soccer analytics (eggels2016expected). This statistic can now be seen on television broadcasts and is even shown as part of the popular video game EA Sports FC (formerly FIFA).
xG quantifies the probability that a shot will result in a goal based on characteristics such as shot location, angle to the goal, shot type, and defensive pressure (spearman2018beyond). Perhaps the earliest use of expected goals was ensum2004applications, which used a logistic regression model to estimate the probability that a shot becomes a goal. macdonald2012expected implements an expected goals model in a different low-scoring sport, ice hockey, where expected goals were used to estimate an adjusted plus-minus model for players in the NHL. Like this work in the NHL, xG in soccer has been shown to be a more predictive metric of future goals scored than actual goals scored (heuer2012perfect & mead2023expected).
lucey2015quality introduced the use of spatio-temporal data to estimate expected goals models. fernandez2019decomposing uses an expected goals model (among others) to decompose the game into a series of decisions and actions by each player.
@@ 1.2 Limitations of Expected Goals
Any limitations of xG are important to recognize as xG is not only used as a metric itself but is a foundational piece to many important research papers in soccer analytics. singh_xT_blog, bransen2018measuring, and statsbomb2021obv use the xG value of actual shots to value on-ball actions in the buildup. This is often referred to as expected threat (xT) and sometimes as on-ball value (OBV).
@@
There have been very few previous attempts to explicitly model the probabilistic nature of shot taking. The work of fernandez2019decomposing and fernandez2021framework decomposes actions into a sequence using spatio-temporal tracking data and models the probability of a pass, ball drive, or shot at the end of a possession. The authors spend little time talking about the expected goals part of their work, focusing on the passing and ball drive components, but do report that they had some model for the probability a shot occurs.
Another work that incorporated some probabilistic component of shot taking is poropudas2021extended, who incorporated a shot decision model into expected threat (xT).
A few attempts at adjusted plus-minus models in soccer using expected goals have also been attempted, similar to macdonald2012expected in hockey. matano2018augmenting attempted to use expected goals in their FIFA rating augmented plus-minus but decided against it due to data limitations. kharrat2020plus and zhang2022cmsac use expected goals to derive various extensions to the augmented plus-minus models proposed by macdonald2012expected.
xG is also typically used in predictive models either alongside or instead of the actual goals in the match, as was the case with FiveThirtyEight’s Soccer Power Index (SPI) (fivethirtyeight_clubsoccerpredictions).
@@
Soccer may also fall victim to the same selection bias that plagues expected points models in American football. brill2025analytics brought attention to this issue and showed that among other issues, the fact that better offensive teams had more plays closer to the end zone affected the certainty of machine learning-based expected points models and those that relied on expected points (such as 4th down models).
@@ 4.1 Baseline Comparison
We benchmark our models against logistic regression baselines trained on four feature sets: (i) ball distance only, (ii) all ball features, (iii) ball + goalkeeper features, and (iv) all features. Each model is evaluated with 5-fold cross-validation on identical splits, and we compare mean out-of-sample log loss. As shown in Table 2, our XGBoost models outperform all logistic baselines – especially on the xS task. Our xG model is conceptually similar to the proprietary models described by hudl_xg_explained.
@@ 4.4 Player Evaluation
Traditional xG-based metrics often suggest that few players consistently “outperform” their expected goals over multiple seasons (goodman2014finishing; statsbomb2017finishing; davis2024biases). In other words, the extent to which a player scores above or below the sum of their xG in one season is only weakly informative about whether they will over- or under-perform in the future.

### v2

@@ 1.1 Expected Goals
Expected Goals (xG) has become the most commonly used metric in modern soccer analytics (Eggels et al., 2016). This statistic can now be seen on television broadcasts and is even shown as part of the popular video game EA Sports FC (formerly FIFA).
xG quantifies the probability that a shot will result in a goal based on characteristics such as shot location, angle to the goal, shot type, and defensive pressure (Spearman, 2018). Perhaps the earliest use of expected goals was Ensum et al. (2004), which used a logistic regression model to estimate the probability that a shot becomes a goal. Macdonald (2012) implements an expected goals model in a different low-scoring sport, ice hockey, where expected goals were used to estimate an adjusted plus-minus model for players in the NHL. Like this work in the NHL, xG in soccer has been shown to be a more predictive metric of future goals scored than actual goals scored (Heuer and Rubner (2012) & Mead et al. (2023)).
Lucey et al. (2015) introduced the use of spatio-temporal data to estimate expected goals models. Fernández et al. (2019) uses an expected goals model (among others) to decompose the game into a series of decisions and actions by each player.
@@ 1.2 Limitations of Expected Goals
Any limitations of xG are important to recognize as xG is not only used as a metric itself but is a foundational piece to many important research papers in soccer analytics. Singh (2018), Bransen and Van Haaren (2018), and Statsbomb (2021) use the xG value of actual shots to value on-ball actions in the buildup. This is often referred to as expected threat (xT) and sometimes as on-ball value (OBV).
@@
There have been very few previous attempts to explicitly model the probabilistic nature of shot taking. The work of Fernández et al. (2019) and Fernández et al. (2021) decomposes actions into a sequence using spatio-temporal tracking data and models the probability of a pass, ball drive, or shot at the end of a possession. The authors spend little time talking about the expected goals part of their work, focusing on the passing and ball drive components, but do report that they had some model for the probability a shot occurs.
Another work that incorporated some probabilistic component of shot taking is Poropudas and Inkilä (2021), who incorporated a shot decision model into expected threat (xT).
A few attempts at adjusted plus-minus models in soccer using expected goals have also been attempted, similar to Macdonald (2012) in hockey. Matano et al. (2018) attempted to use expected goals in their FIFA rating augmented plus-minus but decided against it due to data limitations. Kharrat et al. (2020) and Zhang (2022) use expected goals to derive various extensions to the augmented plus-minus models proposed by Macdonald (2012).
xG is also typically used in predictive models either alongside or instead of the actual goals in the match, as was the case with FiveThirtyEight’s Soccer Power Index (SPI) (FiveThirtyEight, 2020).
@@
Soccer may also fall victim to the same selection bias that plagues expected points models in American football. Brill et al. (2025) brought attention to this issue and showed that among other issues, the fact that better offensive teams had more plays closer to the end zone affected the certainty of machine learning-based expected points models and those that relied on expected points (such as 4th down models).
@@ 4.1 Baseline Comparison
We benchmark our models against logistic regression baselines trained on four feature sets: (i) ball distance only, (ii) all ball features, (iii) ball + goalkeeper features, and (iv) all features. Each model is evaluated with 5-fold cross-validation on identical splits, and we compare mean out-of-sample log loss. As shown in Table 2, our XGBoost models outperform all logistic baselines – especially on the xS task. Our xG model is conceptually similar to the proprietary models described by Hudl (n.d.).
@@ 4.4 Player Evaluation
Traditional xG-based metrics often suggest that few players consistently “outperform” their expected goals over multiple seasons (Goodman, 2014; Kwiatkowski, 2017; Davis and Robberechts, 2024). In other words, the extent to which a player scores above or below the sum of their xG in one season is only weakly informative about whether they will over- or under-perform in the future.